Lost again on a worksheet. The question is:
A vat contains 100 gallons of a 50% brine solution. What is the greatest amount of water that can be added if the salt content cannot drop below 25%.

Answer by ankor@dixie-net.com(22740)

A vat contains 100 gallons of a 50% brine solution. What is the greatest amount of water that can be added if the salt content cannot drop below 25%.
:
Let x = amt of water added for 25% solution
:
100 gal 50% + x gal 0% = 25% of resulting amt
:
.50(100) + 0(x) = .25(100 + x)
50 + 0 = 25 + .25x
50 - 25 = .25x
:
.25x = 25
x = 25/.25
x = 100 gal can be added (max that can be added)
: